# Domain Generalization in-the-Wild: Disentangling Classification from Domain-Aware Representations

### v1

@@ 1 Introduction
Domain generalization (DG) aims to train models that maintain robust performance when encountering out-of-distribution (OOD) data zhou2022domain . A key assumption of DG is that the target domains represent novel data distributions for evaluation. However, this assumption is challenged when evaluating pretrained foundation models like CLIP radford2021learning_CLIP and ALIGN jia2021scaling . These models have been trained on comprehensive web-scale datasets, thus have likely been exposed to most existing domains, contributing to its impressive zero-shot capabilities. Consequently, much research has focused on adapting CLIP through parameter-efficient finetuning zhou2022learning_coop ; zhou2022conditional_cocoop ; gao2024clipadapter ; zhang2022tip , regularization using the original weights wortsman2022robust_wise ; nam2024lipsum ; oh2024towards_carot ; shu2023clipood , and even transductive methods wallingford2023neuralpriming ; martin2024transductive , largely preserving its pretrained knowledge. However, a critical question arises: how would CLIP perform on genuinely unseen domains? A recent study mayilvahanan2024search found that retraining CLIP from scratch using only natural images significantly degrades performance on OOD benchmarks, resulting in performance similar to models trained only on ImageNet. Our results (Figure 1) align with these findings, suggesting current DG evaluations for CLIP may overestimate its true OOD robustness because standard evaluation settings like leave-one-domain-out and existing cross-dataset evaluations may not be sufficiently challenging (Sec. 4.1, 4.2).
@@
Figure 1: Improvement over zeroshot after finetuning on ImageNet (in %). OOD scores are quantified relative to ImageNet (source dataset), illustrating the challenge of DG in-the-wild.
@@
We therefore propose that DG evaluation for foundation models, such as CLIP, should be more challenging, to approximate “domain generalization in-the-wild," where CLIP might encounter diverse and challenging new data in the real-world. We evaluate CLIP on 33 target datasets spanning a diverse range of OODness. To systematically approach evaluation, we quantify a multi-modal OOD score (Sec. 4.2), using ImageNet as both an anchor and a source dataset owing to its inclusion of many classes and concepts. We find that after finetuning on ImageNet, CLIP’s DG performance degrades on datasets with higher OOD scores with respect to ImageNet (Figure 1), consistent with the domain contamination findings mayilvahanan2024search . In addition, to further simulate truly unseen domains, we use an unlearning technique sepahvandselective to make CLIP forget some domains (Sec. 4.3), and find significant performance degradation for existing robust finetuning methods.
@@
Our results (Figure 9), alongside findings on domain contamination mayilvahanan2024search , suggest that for DG in-the-wild, different robust finetuning algorithms are needed for genuinely unseen data. In light of this, we present CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), an end-to-end finetuning method to improve the robustness of CLIP on truly OOD data. A key idea in DG is that learning domain-invariant features is beneficial for robust generalization zhou2022domain ; ganin2016domain_dann . However, naively enforcing domain invariance for a pretrained foundation model could cause catastrophic forgetting of useful features learned from diverse domains during pretraining as the model is forced to make its representations entirely domain-invariant. We hypothesize that to learn effective domain invariance, domain awareness is a prerequisite. This awareness is critical to maintain CLIP’s vast knowledge, which includes generalizable features that support capabilities like zero-shot classification. By enhancing domain awareness, CLIP can also selectively disentangle classification from domain-specific aspects, thereby achieving robust generalization without forgetting valuable information.
@@
We combine the idea of domain awareness and domain invariance by encouraging them simultaneously within CLIP-DCA (Figure 2). Specifically, we encourage domain awareness within CLIP’s image and text encoders, while promoting domain invariance specifically at the final classification layer through disentanglement. Our premise is that while domain awareness is a requirement to maintain pre-existing knowledge, this awareness can be disentangled for domain-invariant classification and robust generalization. To achieve this, we add a new head to the CLIP image encoder, called the domain head, which is trained to understand domains. The original classification head is then disentangled from the domain head, effectively learning domain awareness within its encoders and achieving domain invariance at the classification stage. Additionally, since many datasets lack distinct domains or textual descriptions, and the definition of ‘domain’ is often vague in DG in-the-wild, we address this by using diffusion models to create images of artificial domains and MLLMs to generate descriptions for these artificial domains (Sec. 3.2). Our contributions are summarized as follows:
@@
  • We demonstrate potential limitations in current DG evaluations of foundation models, supported by our results and a recent study. Existing benchmarks may overestimate true OOD robustness, potentially leading finetuning strategies towards in-distribution improvement rather than OOD.
@@
2 Related Work
Domain Generalization. Learning domain-invariant representations has historically been a central idea in domain generalization blanchard2011generalizing ; zhou2022domain . The intuition is that when classifying images from entirely new distributions, learning abstract features common across source domains should provide better robustness for classification in new domains blanchard2011generalizing ; muandet2013domain . Among these, domain-adversarial learning methods have become a relatively standard approach within the DG field due to its conceptual simplicity and effectiveness zhou2022domain . For instance, Domain Adversarial Neural Networks (DANN) ganin2016domain_dann uses an auxiliary domain classifier trained adversarially against the encoder, encouraging the encoder to produce features indistinguishable across source domains. Given the focus of DANN on the central idea of domain invariance, we focus on DANN and its adaptation to CLIP in our analysis. Notably, despite the prevalence of such DG methods, the direct application for CLIP is not well-established and remains underexplored. Naively enforcing domain invariance on foundation models like CLIP, with large pretrained knowledge, risks catastrophic forgetting.
Robust Finetuning of CLIP. The introduction of CLIP marked a significant shift in DG research. The original study radford2021learning_CLIP demonstrated impressive zero-shot classification performance across diverse benchmarks, including OOD datasets. The authors attributed this capability to CLIP learning representations that are less reliant on spurious correlations specific to downstream target datasets, as CLIP was not trained on these specific datasets during its initial pretraining.
The assumption of the inherent OOD robustness in CLIP motivated numerous methods aimed at finetuning CLIP for downstream tasks while enhancing its perceived robustness. A common approach is parameter-efficient finetuning (PEFT) strategies. An early influential study, CoOp zhou2022learning_coop , introduced learnable textual prompts, motivated by observations that manually crafted prompt ensembles improved CLIP’s zero-shot accuracy. Building on this, CoCoOp zhou2022conditional_cocoop made these prompts dynamic by conditioning them on individual image features through a cross-attention mechanism. Similarly, CLIP-Adapter gao2024clipadapter proposed adding lightweight, learnable MLP layers (adapters) to the CLIP encoders, finetuning only these small adapters instead of the entire network. Many more subsequent PEFT methods have also been explored cho2023promptstyler ; chi2024adapting ; lee2025domain ; addepalli2024leveraging ; bai2024soft ; li2022learning ; khattak2025learning ; cheng2024disentangled ; lafon2024gallop .
End-to-end finetuning methods have also been explored, yet many still depend on the original pretrained CLIP weights for regularization or guidance. Wise-FT wortsman2022robust_wise , motivated by observing that standard finetuning often degraded zero-shot OOD performance, ensembles the weights of the finetuned model with the original CLIP weights. CLIP-OOD shu2023clipood used a beta-moving average of the weights during finetuning alongside a regularization term to enhance semantic relationships learned during pretraining. MIRO cha2022domain used mutual information regularization between the finetuning model and the frozen pretrained CLIP model to retain pretrained features.
While many other methods show strong performance on OOD benchmarks, this overview highlights representative approaches, their trends, and assumptions in robust CLIP finetuning. Our work, however, questions whether current evaluation protocols are sufficiently challenging, and suggests the reliance on the pretrained weights may be suboptimal for true OOD generalization, a concern supported by evidence of domain contamination during pretraining mayilvahanan2024search . Consequently, we explore more challenging evaluations and alternative strategies for training CLIP grounded in DG principles.
@@ 3 CLIP-DCA: Disentangling Classification from Enhanced Domain-Aware Representations
The previous sections have highlighted the complexity of evaluating domain generalization in foundational models like CLIP. In light of this, we propose a novel finetuning approach, CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), designed to improve robustness by balancing the trade-off between domain invariance and knowledge retention.
@@ 3.1 Encouraging domain awareness and invariance simultaneously
Enforcing domain invariance in the encoder through conventional domain adversarial learning, for instance, can be harmful. Our experiments show that applying invariance directly leads to worse performance compared to standard finetuning (Figure 8). Forcing the entire model to become domain-invariant can lead to the forgetting of valuable, fine-grained features learned during the pretraining on a large dataset. Conversely, existing CLIP robust finetuning methods discourage divergence from the original pretrained model, and rely on the assumption that CLIP is inherently robust to OOD data. This assumption is challenged by our results (Figure 9) and evidence for domain contamination mayilvahanan2024search .
@@
During training, we use two distinct loss functions for the two types of data we use - the source dataset and generated diffusion images. We use $\ell_{a}$ to refer to agreement loss (the standard CLIP contrastive loss radford2021learning_CLIP or finetuning goyal2023finetune_flyp ). We use $\ell_{d}$ to refer to disentanglement, which we define as the squared sum of the diagonal of the cross-correlation matrix — $\ell_{d}=\sum_{i}((XY^{\mathsf{T}})_{ii})^{2}$, where $X$ and $Y$ are normalized matrices of size [batch, feature]. We also use $P_{T}$ to refer to the projected embeddings of the text encoder, and $P_{H}$ as the projected hidden states of the MLLM.
@@
    • A classification loss (i.e., the standard CLIP contrastive loss goyal2023finetune_flyp ) between the output of the image class head and the text embedding of the class name, $C_{1}:=\ell_{a}(I_{C},P_{T})$.
@@ 3.2 Generating diverse domains
To address this, we construct a small dataset with a diverse number of domains. As illustrated in Figure 4, we prompt a Multimodal Large Language Model (MLLM), specifically LLaVA liu2023visual_llava , to generate ideas of 512 distinct styles for images (e.g. "pixel art"). A text-to-image diffusion model (Stable Diffusion 3 esser2024scaling ) then generates images from these stylistic prompts. We intentionally omit any class labels during image generation to ensure the styles are not biased towards specific classes. We generate 8 images per style, creating a dataset of 4096 images. Finally, the same MLLM generates textual domain descriptions (captions) for each style. We also store the hidden state representations from the MLLM that were used to generate these style descriptions, as these will be used to encourage domain awareness in the text encoder.
@@ 4.1 Evaluating DG in-the-wild performance
While standard domain generalization benchmarks such as DigitsDG zhou2020digitsDG , PACS li2017pacs , Office-Home venkateswara2017officehome , Terra Incognita beery2018terra , FMOW-Wilds and Camelyon-Wilds koh2021fmowwilds , and ImageNet variants russakovsky2015imagenetv1 ; recht2019imagenetv2 ; hendrycks2021imagenetR ; hendrycks2021imagenetA ; wang2019imagenetSketch are widely used, we observe that the different domains within a single benchmark dataset often exhibit greater similarity to each other than to conceptually similar domains across different datasets. To quantify this, we use Spectral-normalized Neural Gaussian Process (SNGP) liu2020simple_sngp to compute the pairwise OOD scores between domains across these benchmarks. We then visualized the pairwise OOD scores with PCA. This analysis reveals significant clustering within individual benchmarks as shown in Figure 5. The clustering within benchmarks, combined with the impressively high zeroshot accuracy, and the success of transductive methods on certain DG datasets wallingford2023neuralpriming ; martin2024transductive , provides strong evidence that the current DG evaluation is not very difficult for CLIP, possibly due to domain contamination.
Consequently, we evaluate using a more challenging cross-dataset setup aiming to simulate DG in-the-wild. We finetune the model on ImageNet-1K deng2009imagenet , and evaluate its generalization capabilities across a wide range of 33 target datasets, including the standard DG benchmarks listed above. A cross-dataset evaluation is significantly more challenging compared to traditional DG setups, as it involves larger visual distribution shifts and also shifts in class labels. This evaluation also aligns with the methodologies of prior studies investigating robust CLIP finetuning zhou2022learning_coop ; zhou2022conditional_cocoop ; gao2024clipadapter ; shu2023clipood , while adding a broader coverage of domains. We use the CLIP ViT-B/32 model for all experiments. Other training details are included in the appendix.
@@ 4.2 Measuring OODness of the target datasets
Given that our DG in-the-wild evaluation includes many target datasets with varying degrees of OODness compared to ImageNet, establishing a quantitative OOD metric is beneficial for a more holistic assessment of OOD robustness. A unique consideration for CLIP is its dual-encoder architecture. To provide a comprehensive score, we utilize OOD measures for both the image and text modalities. For the image encoder, we use SNGP liu2020simple_sngp calibrated on the ImageNet validation data to compute an OOD score for all 33 target datasets. In addition, we use a text-based OOD measure fort2021exploring_clip_text to measure OODness of class labels. This involves calculating classification probabilities on a combined label set of target dataset class names and ImageNet class names using the target domain image embeddings. The text OOD score is the summed probability assigned to the target-specific class names.
We verify that our OOD score shows a strong negative correlation (r=-0.756, p<0.001) with performance on target datasets after finetuning, as shown in Figure 6. Notably, we find that averaging the image and text OOD scores is important for accurately predicting post-finetuning accuracy. Relying solely on the image OOD score (r=-0.099) or the text OOD score (r=-0.608) yields weaker correlations, providing evidence that OOD scores in both modalities are necessary for a comprehensive understanding of OOD challenges in the context of CLIP.
4.3 Approximating unseen data through unlearning
Table 1: Performance of ZS (zero-shot), FT (Regular finetuning on GCC), and Unlearn (Regular finetuning on GCC + unlearning on DomainNet).
@@ 4.3 Approximating unseen data through unlearning
Ideally, evaluating true DG performance would involve using a CLIP model trained without specific target-like domains. Unfortunately, retraining CLIP from scratch while omitting certain data is computationally expensive due to the scale of original training data (millions to billions of images). Public weights for selectively trained models are unavailable.
Given these constraints, we explore an approximate approach inspired by the concept of unlearning to mitigate potential domain contamination. Specifically, we adapt the adversarial learning-based unlearning method sepahvandselective for domain forgetting. We finetune CLIP goyal2023finetune_flyp using a dual objective. First, to retain general knowledge, we train on a 595,000-image subset of the CC3M dataset sharma2018conceptual_cc3m , referred to as GCC, previously used in LLaVA pretraining liu2023visual_llava , serving as a manageable proxy for CLIP’s original training data. Second, to approximate a scenario where domains similar to DomainNet are removed, we apply domain adversarial training ganin2016domain_dann on the DomainNet dataset, which we exclude from our target datasets. We attach a binary classifier to the penultimate layer of the image encoder. During training batches, this classifier is fed representations of random noise (assigned label 0) and images from DomainNet (assigned label 1). The gradient reversal layer ganin2016domain_dann forces the image encoder to learn representations that confuse this classifier, making embeddings of DomainNet images and random noise indistinguishable, thereby encouraging the model to unlearn domain-specific features from DomainNet. The unlearning occurs concurrently with standard training on the GCC dataset to preserve CLIP’s core capabilities.
We intentionally avoid unlearning the target datasets used in our evaluation. This ensures a fairer comparison against existing robust CLIP finetuning methods, many of which discourage heavily diverging from the original pretrained weights. Directly unlearning the target datasets would give the baselines an unfair disadvantage. Instead, unlearning DomainNet serves as a proxy for reducing domain contamination effects. We find that performance on DomainNet and some other datasets drops, as shown in Table 1, while retaining much of the performance on many other datasets.

### v2

@@ 1 Introduction
Figure 1: Improvement over zeroshot after finetuning on ImageNet (in %). Each dot represents a target dataset. OOD scores are quantified relative to ImageNet (source dataset), illustrating the challenge of DG in-the-wild.
Domain generalization (DG) aims to train models that maintain robust performance when encountering out-of-distribution (OOD) data (Zhou et al., 2022a). A key assumption of DG is that the target domains represent novel data distributions for evaluation. However, this assumption is challenged when evaluating pretrained foundation models like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These models have been trained on comprehensive web-scale datasets, thus have likely been exposed to most existing domains, contributing to its impressive zero-shot capabilities. Consequently, much research has focused on adapting CLIP through parameter-efficient finetuning (Zhou et al., 2022c; b; Gao et al., 2024; Zhang et al., 2022), regularization using the original weights (Wortsman et al., 2022; Nam et al., 2024; Oh et al., 2024; Shu et al., 2023), and even transductive methods (Wallingford et al., 2023; Martin et al., 2024), largely preserving its pretrained knowledge. However, this reliance on pretrained knowledge is predicated on an assumption of true OOD robustness that is now being challenged. Recent studies show that evaluating foundation models for DG is often compromised by data leakage from web-scale pre-training (Teterwak et al., 2024; Yu et al., 2024). Teterwak et al. (2024) addresses this by analyzing generalization from learned to unlearned samples within the pre-training data, while Yu et al. (2024) proposes training models from scratch to avoid contamination entirely. These studies, along with findings that retraining CLIP on cleaner data degrades OOD performance, suggest that current DG evaluations may overestimate true OOD robustness (Mayilvahanan et al., 2024). While these studies provide critical insight, they do not offer a tractable setting to test the generalization performance of CLIP after specific, contaminating knowledge has been selectively removed.
@@
To address this gap, we propose that DG evaluation for foundation models, such as CLIP, should be more challenging, to approximate “domain generalization in-the-wild," where CLIP might encounter diverse and challenging new data in the real-world. We evaluate CLIP on 33 target datasets spanning a diverse range of OODness. To systematically approach evaluation, we quantify a multi-modal OOD score (Sec. 3.2), using ImageNet as both an anchor and a source dataset owing to its inclusion of many classes and concepts. We find that after finetuning on ImageNet, CLIP’s DG performance degrades on datasets with higher OOD scores with respect to ImageNet (Figure 1), consistent with the domain contamination findings (Mayilvahanan et al., 2024). In addition, to further simulate truly unseen domains, we use an unlearning technique (Sepahvand et al., 2025) to make CLIP forget some domains (Sec. 3.3), and find significant performance degradation for existing robust finetuning methods.
@@
Our results (Figure 9), alongside findings on domain contamination (Mayilvahanan et al., 2024), suggest that for DG in-the-wild, different robust finetuning algorithms are needed for genuinely unseen data. In light of this, we present CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), an end-to-end finetuning method to improve the robustness of CLIP on truly OOD data. A key idea in DG is that learning domain-invariant features is beneficial for robust generalization (Zhou et al., 2022a; Ganin et al., 2016). However, naively enforcing domain invariance for a pretrained foundation model could cause catastrophic forgetting of useful features learned from diverse domains during pretraining as the model is forced to make its representations entirely domain-invariant. We hypothesize that to learn effective domain invariance, domain awareness is a prerequisite. This awareness is critical to maintain CLIP’s vast knowledge, which includes generalizable features that support capabilities like zero-shot classification. By enhancing domain awareness, CLIP can also selectively disentangle classification from domain-specific aspects, thereby achieving robust generalization without forgetting valuable information.
@@
We combine the idea of domain awareness and domain invariance by encouraging them simultaneously within CLIP-DCA (Figure 2). Specifically, we encourage domain awareness within CLIP’s image and text encoders, while promoting domain invariance specifically at the final classification layer through disentanglement. Our premise is that while domain awareness is a requirement to maintain pre-existing knowledge, this awareness can be disentangled for domain-invariant classification and robust generalization. To achieve this, we add a new head to the CLIP image encoder, called the domain head, which is trained to understand domains. The original classification head is then disentangled from the domain head, effectively learning domain awareness within its encoders and achieving domain invariance at the classification stage. Additionally, since many datasets lack distinct domains or textual descriptions, and the definition of ‘domain’ is often vague in DG in-the-wild, we address this by using diffusion models to create images of artificial domains and Multimodal LLMs (MLLMs) to generate descriptions for these artificial domains (Sec. 2.2). Our contributions are summarized as follows:
@@
  • We demonstrate potential limitations in current DG evaluations of foundation models, supported by our results and recent studies. Existing benchmarks may overestimate true OOD robustness, potentially leading finetuning strategies towards in-distribution improvement rather than OOD.
@@
Related Work. A comprehensive review is in Appendix A. Domain generalization (DG) has traditionally focused on learning domain-invariant representations (Ganin et al., 2016; Zhou et al., 2022a), but naively applying these methods to foundation models like CLIP can cause catastrophic forgetting. Consequently, most robust CLIP finetuning methods aim to preserve pretrained knowledge through parameter-efficient finetuning (PEFT) (Zhou et al., 2022c; Gao et al., 2024) or regularization towards the original weights (Wortsman et al., 2022; Shu et al., 2023). However, this reliance on pretrained knowledge is being questioned by recent findings of domain contamination in web-scale datasets (Teterwak et al., 2024; Yu et al., 2024; Mayilvahanan et al., 2024), which suggest current evaluations may overestimate true OOD robustness. Our work addresses this by proposing a more challenging evaluation framework and a method that learns targeted invariance without sacrificing pretrained knowledge.
@@ 2 CLIP-DCA: Disentangling Classification from Enhanced Domain-Aware Representations
To address the challenges of DG in-the-wild, we introduce CLIP-DCA (Disentangling Classification from enhanced domain Aware representations), a finetuning method designed to improve robustness on genuinely unseen data.
@@ 2.1 Encouraging domain awareness and invariance simultaneously
Enforcing domain invariance in the encoder through conventional domain adversarial learning, for instance, can be harmful. Our experiments show that applying invariance directly leads to worse performance compared to standard finetuning (Figure 8). Forcing the entire model to become domain-invariant can lead to the forgetting of valuable, fine-grained features learned during the pretraining on a large dataset. Conversely, existing CLIP robust finetuning methods discourage divergence from the original pretrained model, and rely on the assumption that CLIP is inherently robust to OOD data. This assumption is challenged by our results (Figure 9) and evidence for domain contamination (Teterwak et al., 2024; Yu et al., 2024; Mayilvahanan et al., 2024).
@@
During training, we use two distinct loss functions for the two types of data we use - the source dataset and generated diffusion images. We use $\ell_{a}$ to refer to agreement loss (the standard CLIP contrastive loss (Radford et al., 2021) or finetuning (Goyal et al., 2023)). We use $\ell_{d}$ to refer to disentanglement, which enforces statistical independence between two sets of representations. Inspired by the simplicity of self-supervised methods (Zbontar et al., 2021; Bardes et al., 2021), we achieve this by minimizing the correlation between the class and domain embeddings. The loss is formulated as the squared sum of the diagonal of the cross-correlation matrix between the batch-normalized class embeddings and domain embeddings. This penalizes any shared information, encouraging the class head to find predictive representations that are independent of features useful for domain prediction.
The role of the disentanglement loss is to enforce this separation. The underlying assumption is that if the two representations are truly disentangled, the features from the class head for a given sample should be statistically independent from the features learned by the domain head for that same sample. By minimizing the correlation between the class and domain embeddings, this loss encourages the class head to find representations that are predictive of the class label without using features that are also useful for predicting the domain. Conversely, it encourages the domain head to focus only on domain-specific information, as any shared information with the class head is penalized.
@@
    • A classification loss (i.e., the standard CLIP contrastive loss (Goyal et al., 2023)) between the output of the image class head and the text embedding of the class name, $C_{1}:=\ell_{a}(I_{C},P_{T})$.
@@
For a detailed implementation, pseudocode for the main training loop is provided in Appendix B.
@@ 2.2 Generating diverse domains
To address this, we construct a small dataset with a diverse number of domains. As illustrated in Figure 4, we prompt a MLLM, specifically LLaVA (Liu et al., 2023), to generate ideas of 512 distinct styles for images (e.g. "pixel art"). The complete list of styles is available in LABEL:tab:synthetic_styles_list in the Appendix. A text-to-image diffusion model (Stable Diffusion 3 (Esser et al., 2024)) then generates images from these stylistic prompts. We intentionally omit any class labels during image generation to ensure the styles are not biased towards specific classes. We generate 8 images per style, creating a dataset of 4096 images. Finally, the same MLLM generates textual domain descriptions (captions) for each style. We also store the hidden state representations from the MLLM that were used to generate these style descriptions, as these will be used to encourage domain awareness in the text encoder. The exact prompts used for style and description generation are detailed in Appendix F.
@@ 3.1 Evaluating DG in-the-wild performance
We first analyze standard Domain Generalization (DG) benchmarks and find their domains are not well-separated. Using a Spectral-normalized Neural Gaussian Process (SNGP) (Liu et al., 2020) to compute pairwise OOD scores, we observe strong intra-benchmark clustering, as visualized in Figure˜5. This clustering, along with CLIP’s high zero-shot accuracy and the success of transductive methods on these datasets (Wallingford et al., 2023; Martin et al., 2024), suggests that current DG evaluations are not sufficiently challenging for large-scale models, possibly due to pre-training data contamination.
To address this, we finetune CLIP on ImageNet-1K (Deng et al., 2009) and evaluate its generalization capabilities across a more diverse benchmark of 33 target datasets spanning standard DG benchmarks and other challenging classification tasks (full list provided in LABEL:tab:datasets_list_stats). A cross-dataset evaluation is significantly more challenging compared to traditional DG setups, as it involves larger visual distribution shifts and also shifts in class labels. This evaluation also aligns with the methodologies of prior studies investigating robust CLIP finetuning (Zhou et al., 2022c; b; Gao et al., 2024; Shu et al., 2023), while adding a broader coverage of domains. We use the CLIP ViT-B/32 model for all experiments. Further implementation details, including optimizer settings and specific hyperparameters for our method, are provided in Appendix C and D.
@@ 3.2 Measuring OODness of the target datasets
Given that our DG in-the-wild evaluation includes many target datasets with varying degrees of OODness compared to ImageNet, establishing a quantitative OOD metric is beneficial for a more holistic assessment of OOD robustness. A unique consideration for CLIP is its dual-encoder architecture. To provide a comprehensive score, we utilize OOD measures for both the image and text modalities. For the image encoder, we use SNGP (Liu et al., 2020) calibrated on the ImageNet validation data to compute an OOD score for all 33 target datasets. In addition, we use a text-based OOD measure (Fort et al., 2021) to measure OODness of class labels. This involves calculating classification probabilities on a combined label set of target dataset class names and ImageNet class names using the target domain image embeddings. The text OOD score is the summed probability assigned to the target-specific class names.
We verify that our OOD score shows a strong negative correlation (r=-0.756, p<0.001) with performance on target datasets after finetuning, as shown in Figure 6. Notably, we find that averaging the image and text OOD scores is important for accurately predicting post-finetuning accuracy. Relying solely on the image OOD score (r=-0.099) or the text OOD score (r=-0.608) yields weaker correlations, providing evidence that OOD scores in both modalities are necessary for a comprehensive understanding of OOD challenges in the context of CLIP. A detailed breakdown of the OOD scores for all 33 target datasets is provided in Appendix J.
Table 1: Unlearning effectiveness. ZS: Original zero-shot performance. FT: Baseline fine-tuning on the GCC retention set. Unlearn: Full unlearning combining retention on GCC with adversarial unlearning on DomainNet.
@@
3.3 Simulating Unseen Domains via Unlearning
Retraining a foundation model like CLIP from scratch to omitting specific domains is computationally prohibitive. To overcome this, we use an unlearning method as a proxy to approximate a model that is not contaminated with domains relevant to our evaluation. This controlled experiment allows us to answer a critical research question: How do robust fine-tuning methods perform on genuinely unseen domains? Our results expose weaknesses in existing approaches that rely on pretrained weights.
Specifically, we adapt the adversarial learning-based unlearning method (Sepahvand et al., 2025) for domain forgetting. We finetune CLIP (Goyal et al., 2023) using a dual objective. First, to retain general knowledge, we train on a 595,000-image subset of the CC3M dataset (Sharma et al., 2018), referred to as GCC, previously used in LLaVA pretraining (Liu et al., 2023), serving as a manageable proxy for CLIP’s original training data. Second, to approximate a scenario where domains similar to DomainNet are removed, we apply domain adversarial training (Ganin et al., 2016) on the DomainNet dataset, which we exclude from our target datasets. We attach a binary classifier to the penultimate layer of the image encoder. During training batches, this classifier is fed representations of random noise (assigned label 0) and images from DomainNet (assigned label 1). The gradient reversal layer (Ganin et al., 2016) forces the image encoder to learn representations that confuse this classifier, making embeddings of DomainNet images and random noise indistinguishable, thereby encouraging the model to unlearn domain-specific features from DomainNet. The unlearning occurs concurrently with standard training on the GCC dataset to preserve CLIP’s core capabilities. A pseudocode of the unlearning process is in Appendix B.
We deliberately unlearn on DomainNet, a dataset we do not use for final evaluation. Unlearning our target evaluation datasets directly would unfairly penalize baseline methods. Many methods are designed to regularize against large deviations from the original pretrained weights. By using DomainNet as a proxy for domain contamination, we ensure a fairer comparison. The effectiveness of our unlearning is confirmed by a performance drop on DomainNet while performance on many other datasets is largely retained (Table 1).
This experimental setup is distinct from other recent proposals. While Teterwak et al. (2024) separate samples based on whether they were learned during pre-training, their focus is on generalization from well-learned to seen-but-unlearned concepts. In contrast, Yu et al. (2024) evaluate domain generalization by training models from scratch without web-scale pre-training. Our approach is a unique and practical middle ground. We measure the performance of a model that benefits from web-scale pre-training but has had specific domain knowledge removed. This allows us to more directly isolate the effect of domain contamination on robust fine-tuning.